# Thermodynamic nonequilibrium effects in three-dimensional high-speed compressible flows: Multiscale modeling and simulation via the discrete Boltzmann method

### v1

@@ II.3 Derivation and extraction of thermodynamic nonequilibrium effects
$\partial_{t}(\rho e)+\bm{\nabla}\cdot[(e+P)\mathbf{u}+\bm{\Delta}_{2}^{*}\cdot\mathbf{u}+\bm{\Delta}_{31}^{*}]=0,$ (27)
where $e=c_{v}T+u^{2}/2$ denotes the energy density, and $c_{v}=(n+3)R/2$ is the specific heat at constant volume.
From the perspective of kinetic macroscopic modeling, the physical functions of DBM correspond to the extended hydrodynamic equations, which encompass not only the evolution of conserved moments, such as mass, momentum, and energy, but also the evolution of closely related nonconserved moments. The evolution of low-order conserved kinetic moments is fundamental to hydrodynamics, while high-order nonconserved moments become more important with increased discretization and stronger nonequilibrium effects.
@@
$\bm{\Delta}_{mn}=\mathbf{M}_{mn}(f_{i}-f_{i}^{(0)})=\sum_{i}\left(\frac{1}{2}\right)^{1-\delta_{mn}}$ (28)
@@
$\bm{\Delta}_{mn}^{*}=\mathbf{M}_{mn}^{*}(f_{i}-f_{i}^{(0)})=\sum_{i}\left(\frac{1}{2}\right)^{1-\delta_{mn}}$ (29)
$(f_{i}-f_{i}^{(0)})\underbrace{\mathbf{v}_{i}^{*}\mathbf{v}_{i}^{*}\cdots\mathbf{v}_{i}^{*}}_{n}({v}_{i}^{*2}+\eta_{i}^{*2})^{\frac{m-n}{2}},$
where $\delta_{mn}$ is the Kronecker delta function, and $\mathbf{v}_{i}^{*}=\mathbf{v}_{i}-\mathbf{u}$, $\bm{\Delta}_{mn}^{*}$ represents the contraction of an $m$-th order tensor to an $n$-th order one. When $m=n$, $\bm{{\Delta}}_{mn}^{*}$ is abbreviated as $\bm{{\Delta}}_{m}^{*}$. If $\mathbf{v}_{i}^{*}$ is replaced by $\mathbf{v}_{i}$, the central moment $\bm{\Delta}_{mn}^{*}$ transforms into the non-central moment $\bm{\Delta}_{mn}$. The central moment $\bm{\Delta}_{mn}^{*}$ captures the thermal fluctuations of microscopic particles relative to $\mathbf{u}$, describing only TNEs. In contrast, $\bm{\Delta_{mn}}$ includes both flow velocity and thermal fluctuations, capturing both HNE and TNE effects. These two descriptions are interconnected and complementary, providing unique perspectives for describing or measuring the system’s deviation from thermodynamic equilibrium. Additionally, the independent components of $\bm{\Delta_{mn}}$ and $\bm{\Delta_{mn}}^{*}$ can form the basis for constructing a phase space, offering a geometric framework for representing the states and behavior of complex systems. The use of nonconserved kinetic moments of $(f_{i}-f_{i}^{(0)})$ to detect and describe deviation from thermodynamic equilibrium, along with their associated effects, represents a key feature of DBM in complex physical field analysis.
Two of the most critical nonequilibrium quantities in fluid modeling are the viscous stress ${\bm{\Delta}}{2}^{*}$ and heat flux ${\bm{\Delta}}{31}^{*}$, for which we derive analytical expressions using the CE multiscale expansion. Other relevant nonequilibrium quantities can be derived similarly.
Taking the velocity moments with the collision-invariant vector $\left(1,\mathbf{v}i,(v_{i}^{2}+\eta_{i}^{2})/2\right)$ on both sides of Eq. (9), we obtain the following relations between the temporal derivative $\partial{t_{1}}$ and the spatial derivative $\bm{\nabla}_{1}$:
@@
$\bm{\Delta}_{2}^{*(1)}=\sum_{i}f^{(1)}\mathbf{v}_{i}^{*}\mathbf{v}_{i}^{*}=-\mu\left[\bm{\nabla}\mathbf{u}+\left(\bm{\nabla}\mathbf{u}\right)^{T}-\frac{2}{n+3}\mathbf{I}\bm{\nabla}\cdot\mathbf{u}\right]=-\sigma_{\text{NS}},$ (33)
$\Delta_{31}^{*(1)}=\sum_{i}\frac{1}{2}f^{(1)}\mathbf{v}_{i}^{*}\cdot\mathbf{v}_{i}^{*}\mathbf{v}_{i}^{*}=-\kappa\bm{\nabla}T=-\mathbf{j}_{q,\text{NS}},$ (34)
@@
$\partial_{t_{2}}\big{[}\frac{(n+3)}{2}\rho T+\frac{1}{2}\rho u^{2}\big{]}=-\bm{\nabla}_{1}\cdot\bm{\Delta}_{31}^{(1)}.$ (37)
@@
$\bm{\Delta}_{2}^{*(2)}=\sum_{i}f_{i}^{(2)}\mathbf{v}_{i}^{*}\mathbf{v}_{i}^{*}=-(\sigma_{\text{Burnett }}-\sigma_{\text{NS}}),\$ (38)
$\bm{\Delta}_{31}^{*(2)}=\sum_{i}f_{i}^{(2)}\frac{v_{i}^{*2}+\eta_{i}^{*2}}{2}\mathbf{v}_{i}^{*}=-(\mathbf{j}_{q,\text{Burnett}}-\mathbf{j}_{q,\text{NS}}).$ (39)
These equations provide explicit expressions for the second-order constitutive relations. Here, $(\sigma_{\text{Burnett}}-\sigma_{\text{NS}})$ represents the components where the NS equations fail to describe the system within the range of applicability of the Burnett equations. Therefore, the second-order constitutive relations are given by:
@@
$\bm{\Delta}_{31}^{*}=\bm{\Delta}_{31}^{*(1)}+\bm{\Delta}_{31}^{*(2)}.$ (41)
Meanwhile, we point out that using a single relaxation time in the collision term fixes the Prandtl number at $\Pr=1$. To overcome this limitation, an external forcing term, $I_{i}=[ART+B(\mathbf{v}_{i}-\mathbf{u})^{2}]f_{i}^{(0)}$, is added to the right-hand side of Eq. (4) to modify the BGK collision operator Gan et al. (2011a). Here, $A=-2B$ and $B=\frac{1}{2\rho T^{2}}\bm{\nabla}\cdot\left[\frac{5+n}{2}\rho Tq\bm{\nabla}T\right]$. Consequently, the heat conductivity becomes $\kappa=c_{p}P(\tau+q)$, and the Prandtl number is updated as $\Pr=\tau/(\tau+q)$. This method is straightforward and eliminates the need for $f^{(0)}$ to satisfy additional kinetic moments.
@@ III Numerical tests and analysis
In this section, several typical test cases, ranging from 1D to 3D, are used to evaluate the model’s capability to capture large-scale flow structures. To ensure high accuracy, stability, and computational efficiency, we employ the following numerical schemes: the time derivative is discretized using a second-order implicit-explicit Runge-Kutta scheme Ascher, Ruuth, and Spiteri (1997), and the spatial derivative is approximated using either a fifth-order weighted essentially non-oscillatory (5th WENO) scheme Liu, Osher, and Chan (1994) or a second-order non-oscillatory, parameter-free dissipative (2nd NND) scheme. The discrete Boltzmann equation, particle velocity, and hydrodynamic quantities are nondimensionalized using appropriate reference variables Gan et al. (2011b). The three independent reference variables are the characteristic flow length scale $L_{0}$, the reference density $\rho_{0}$, and the reference temperature $T_{0}$. Additional derived reference variables are defined as $u_{0}=\sqrt{RT_{0}}$, $t_{0}=L_{0}/u_{0}$, and $P_{0}=\rho_{0}RT_{0}$, where $R=287~{}\mathrm{J/(kg~{}K)}$ is the specific gas constant. The simulations assume the fluid is air under standard conditions. Accordingly, the reference values are set to $\rho_{0}=1.165~{}\mathrm{kg/m^{3}}$, $T_{0}=303~{}\mathrm{K}$, and $u_{0}=\sqrt{RT_{0}}\approx 294.892~{}\mathrm{m/s}$.
@@ III.1.1 Einfeldt shock tube
$\left\{\begin{array}[]{l}\left(\rho,T,u_{x},u_{y}\right)|_{L}=\left(1.0,1.0,0.75,0.0\right),\\ \left(\rho,T,u_{x},u_{y}\right)|_{R}=\left(0.125,0.8,0.0,0.0\right).\end{array}\right.$ (42)
@@ III.1.2 Modified Sod shock tube
$\left\{\begin{array}[]{l}\left(\rho,T,u_{x},u_{y}\right)|_{L}=\left(1.0,1.0,0.3,0.0\right),\\ \left(\rho,T,u_{x},u_{y}\right)|_{R}=\left(0.125,0.25,-0.1,0.0\right).\end{array}\right.$ (47)
@@ III.1.3 Lax shock tube
$\left\{\begin{array}[]{l}\left(\rho,T,u_{x},u_{y}\right)|_{L}=\left(0.445,7.928,0.698,0.0\right),\\ \left(\rho,T,u_{x},u_{y}\right)|_{R}=\left(0.50,1.142,0.0,0.0\right).\end{array}\right.$ (48)
@@ III.1.4 Collision of two strong shocks
$\left\{\begin{array}[]{l}\left(\rho,T,u_{x},u_{y}\right)|_{L}=\left(5.99924,76.8254,19.5975,0.0\right),\\ \left(\rho,T,u_{x},u_{y}\right)|_{R}=\left(5.99242,7.69222,-6.19633,0.0\right).\end{array}\right.$ (49)
The numerical parameters for this problem are $\Delta x=\Delta y=\Delta z=4\times 10^{-3}$, $\Delta t=5\times 10^{-5}$, $\tau=5\times 10^{-5}$, $n=0$, $c=-13$, and $\eta_{0}=34$, while the remaining parameters remain unchanged. The numerical solutions for density, pressure, velocity, and temperature distributions closely align with the analytical solutions. The model handles strong shock waves with remarkable stability and accuracy, exhibiting minimal numerical dissipation. It is worth noting that slight diffusion is observed in the steep shock wave regions. This diffusion is attributed to the DBM’s inherent dissipative mechanisms, such as viscous stress and heat conduction, enabling more effective representation of meso-scale structures like shock interfaces. This confirms the model’s robustness and stability in handling strong nonlinear compressible flows.
@@
This study analyzes six selected configurations, with their initial conditions provide in Table 1. The symbols $(P_{i},\rho_{i},u_{xi},u_{yi})$ represent the physical quantities in the $i$th quadrant. The simulation employs a grid of $N_{x}\times N_{y}\times N_{z}=200\times 200\times 4$, applying Dirichlet boundary conditions in the $x$ direction and periodic conditions in the $y$ and $z$ directions. The grid spacing and time step are set as $\Delta x=\Delta y=\Delta z=10^{-3}$, $\Delta t=10^{-4}$, and the relaxation time as $\tau=5\times 10^{-5}$.

### v2

@@ II.3 Derivation and extraction of thermodynamic nonequilibrium effects
$\partial_{t}(\rho e)+\bm{\nabla}\cdot[(\rho e+P)\mathbf{u}+\bm{\Delta}_{2}^{*}\cdot\mathbf{u}+\bm{\Delta}_{3,1}^{*}]=0,$ (27)
where $P=\rho RT$ is the ideal gas equation of state, $e=c_{v}T+u^{2}/2$ denotes the energy density, and $c_{v}=(n+3)R/2$ is the specific heat at constant volume.
The CE multiscale analysis provides the theoretical foundation for such kinetic models. DBM extends beyond the generalized hydrodynamic equations from CE analysis in capturing system properties. From the perspective of kinetic macroscopic modeling, the physical functions of DBM correspond to the extended hydrodynamic equations, which encompass not only the evolution of conserved kinetic moments, such as mass, momentum, and energy, but also the evolution of closely related nonconserved kinetic moments. For example, evolution equations of $\bm{\Delta}_{m}^{*}$ and $\bm{\Delta}_{m,n}^{*}$ areGan et al. (2022, 2025)
$\partial_{t}\bm{\Delta}_{m}^{*}+\partial_{t}\mathbf{M}_{m}^{*}\left(f_{i}^{(0)}\right)+\bm{\nabla}\cdot\left[\mathbf{M}_{m+1}^{*}\left(f_{i}^{(0)}\right)\right.$ (28)
$\left.+\mathbf{M}_{m}^{*}\left(f_{i}^{(0)}\right)\mathbf{u}+\bm{\Delta}_{m+1}^{*}+\bm{\Delta}_{m}^{*}\mathbf{u}\right]=-\frac{1}{\tau}\bm{\Delta}_{m}^{*},$
$\partial_{t}\bm{\Delta}_{m,n}^{*}+\partial_{t}\mathbf{M}_{m,n}^{*}\left(f_{i}^{(0)}\right)+\bm{\nabla}\cdot\left[\mathbf{M}_{m+1,n+1}^{*}\left(f_{i}^{(0)}\right)\right.$ (29)
$\left.+\mathbf{M}_{m,n}^{*}\left(f_{i}^{(0)}\right)\mathbf{u}+\bm{\Delta}_{m+1,n+1}^{*}+\bm{\Delta}_{m,n}^{*}\mathbf{u}\right]=-\frac{1}{\tau}\bm{\Delta}_{m,n}^{*}.$
The evolution of low-order conserved kinetic moments is fundamental to hydrodynamics, while high-order nonconserved kinetic moments become more important with increased discretization and stronger nonequilibrium effects.
@@
$\bm{\Delta}_{m,n}=\mathbf{M}_{m,n}(f_{i}-f_{i}^{(0)})=\sum_{i}\left(\frac{1}{2}\right)^{1-\delta_{mn}}$ (30)
@@
$\bm{\Delta}_{m,n}^{*}=\mathbf{M}_{m,n}^{*}(f_{i}-f_{i}^{(0)})=\sum_{i}\left(\frac{1}{2}\right)^{1-\delta_{mn}}$ (31)
$(f_{i}-f_{i}^{(0)})\underbrace{\mathbf{v}_{i}^{*}\mathbf{v}_{i}^{*}\cdots\mathbf{v}_{i}^{*}}_{n}({v}_{i}^{*2}+\eta_{i}^{2})^{\frac{m-n}{2}},$
where $\delta_{mn}$ is the Kronecker delta function, and $\mathbf{v}_{i}^{*}=\mathbf{v}_{i}-\mathbf{u}$, $\bm{\Delta}_{mn}^{*}$ represents the contraction of an $m$th order tensor to an $n$th order one. When $m=n$, $\bm{{\Delta}}_{m,n}^{*}$ is abbreviated as $\bm{{\Delta}}_{m}^{*}$. If $\mathbf{v}_{i}^{*}$ is replaced by $\mathbf{v}_{i}$, the central moment $\bm{\Delta}_{m,n}^{*}$ transforms into the non-central moment $\bm{\Delta}_{m,n}$. The central moment $\bm{\Delta}_{m,n}^{*}$ captures the thermal fluctuations of microscopic particles relative to $\mathbf{u}$, describing only TNEs. In contrast, $\bm{\Delta_{m,n}}$ includes both flow velocity and thermal fluctuations, capturing both HNE and TNE effects. These two descriptions are interconnected and complementary, providing unique perspectives for describing or measuring the system’s deviation from thermodynamic equilibrium. Additionally, the independent components of $\bm{\Delta_{m,n}}$ and $\bm{\Delta_{m,n}}^{*}$ can form the basis for constructing a phase space, offering a geometric framework for representing the states and behavior of complex systems. The use of nonconserved kinetic moments of $(f_{i}-f_{i}^{(0)})$ to detect and describe deviation from thermodynamic equilibrium, along with their associated effects, represents a key feature of DBM in complex physical field analysis.
Two of the most critical nonequilibrium quantities in fluid modeling are the viscous stress ${\bm{\Delta}}_{2}^{*}$ and heat flux ${\bm{\Delta}}_{3,1}^{*}$, for which we derive analytical expressions using the CE multiscale expansion.
Taking the velocity moments with the collision-invariant vector $\left(1,\mathbf{v}_{i},(v_{i}^{2}+\eta_{i}^{2})/2\right)$ on both sides of Eq. (9), we obtain the following relations between the temporal derivative $\partial_{t_{1}}$ and the spatial derivative $\bm{\nabla}_{1}$:
@@
$\bm{\Delta}_{2}^{*(1)}=\sum_{i}f_{i}^{(1)}\mathbf{v}_{i}^{*}\mathbf{v}_{i}^{*}=-\mu\left[\bm{\nabla}\mathbf{u}+\left(\bm{\nabla}\mathbf{u}\right)^{T}-\frac{2}{n+3}\mathbf{I}\bm{\nabla}\cdot\mathbf{u}\right]=-\bm{\sigma}_{\text{NS}},$ (35)
$\bm{\Delta}_{3,1}^{*(1)}=\sum_{i}\frac{1}{2}f_{i}^{(1)}\frac{v_{i}^{*2}+\eta_{i}^{2}}{2}\mathbf{v}_{i}^{*}=-\kappa\bm{\nabla}T=-\mathbf{j}_{q,\text{NS}},$ (36)
@@
$\partial_{t_{2}}\big{[}\frac{(n+3)}{2}\rho T+\frac{1}{2}\rho u^{2}\big{]}=-\bm{\nabla}_{1}\cdot\bm{\Delta}_{3,1}^{(1)}.$ (39)
@@
$\bm{\Delta}_{2}^{*(2)}=\sum_{i}f_{i}^{(2)}\mathbf{v}_{i}^{*}\mathbf{v}_{i}^{*}=-(\bm{\sigma}_{\text{Burnett }}-\bm{\sigma}_{\text{NS}}),\$ (40)
$\bm{\Delta}_{3,1}^{*(2)}=\sum_{i}f_{i}^{(2)}\frac{v_{i}^{*2}+\eta_{i}^{2}}{2}\mathbf{v}_{i}^{*}=-(\mathbf{j}_{q,\text{Burnett}}-\mathbf{j}_{q,\text{NS}}).$ (41)
These equations provide explicit expressions for the second-order constitutive relations. These equations can not only improve traditional macroscopic fluid modeling but also guide cross-scale regulation of nonequilibrium effects. Here, $(\bm{\sigma}_{\text{Burnett}}-\bm{\sigma}_{\text{NS}})$ represents the components where the NS equations fail to describe the system within the range of applicability of the Burnett equations. Therefore, the second-order constitutive relations are given by:
@@
$\bm{\Delta}_{3,1}^{*}=\bm{\Delta}_{3,1}^{*(1)}+\bm{\Delta}_{3,1}^{*(2)}.$ (43)
Other higher-order thermodynamic nonequilibrium quantities can also be analytically derived similarly, for example, $\Delta_{3xxx}^{*(1)}=-3\tau\rho R^{2}T{\partial x}T$, $\Delta_{3xxx}^{*(2)}=6n_{3}^{-1}\tau^{2}\rho R^{2}T\Big{[}n_{1}T\frac{\partial^{2}}{\partial x^{2}}u_{x}-2T\frac{\partial^{2}}{\partial x\partial y}u_{y}-2T\frac{\partial^{2}}{\partial x\partial z}u_{z}+\partial_{x}T\big{(}(3n+5)\partial_{x}u_{x}-4\partial_{y}u_{y}-4\partial_{z}u_{z}\big{)}+n_{3}\big{(}\partial_{y}T\partial_{y}u_{x}+\partial_{z}T\partial_{z}u_{x}\big{)}\Big{]}$.
Meanwhile, we point out that using a single relaxation time in the collision term fixes the Prandtl number at $\Pr=1$. To overcome this limitation, an external forcing term, $I_{i}=[ART+B(\mathbf{v}_{i}-\mathbf{u})^{2}]f_{i}^{(0)}$, is added to the right-hand side of Eq. (4) to modify the BGK collision operator Gan et al. (2011a). Here, $A=-3B$ and $B=\frac{1}{3\rho T^{2}}\bm{\nabla}\cdot\left[\frac{5+n}{2}\rho Tq\bm{\nabla}T\right]$. Consequently, the heat conductivity becomes $\kappa=c_{p}P(\tau+q)$, and the Prandtl number is updated as $\Pr=\tau/(\tau+q)$. This method is straightforward and eliminates the need for $f^{(0)}$ to satisfy additional kinetic moments. We also point out that, this approach is based on the equilibrium distribution function. It allows for the introduction of HNEs, but does not introduce additional TNEs. To further improve the model, advanced schemes such as the Shakhov model, elliptic statistical model, multi-relaxation time, or two-relaxation time approaches can be considered.
@@ III Numerical tests and analysis
In this section, several typical test cases, ranging from 1D to 3D, are used to evaluate the model’s capability to capture large-scale flow structures. To ensure high accuracy, stability, and computational efficiency, we employ the following numerical schemes: the time derivative is discretized using a second-order implicit-explicit Runge-Kutta scheme Ascher, Ruuth, and Spiteri (1997), and the spatial derivative is approximated using either a fifth-order weighted essentially non-oscillatory (fifth WENO) scheme Liu, Osher, and Chan (1994) or a second-order non-oscillatory, parameter-free dissipative (second NND) scheme Zhang (1988). As the extent of nonequilibrium effects increases, higher-order numerical schemes become necessary to maintain numerical solution accuracy and stability. The discrete Boltzmann equation, particle velocity, and hydrodynamic quantities are nondimensionalized using appropriate reference variables Gan et al. (2011b). The three independent reference variables are the characteristic flow length scale $L_{0}$, the reference density $\rho_{0}$, and the reference temperature $T_{0}$. Consequently, dimensionless variables are $\hat{r}_{\alpha}=\frac{r_{\alpha}}{L_{0}}$, $(\hat{v}_{\alpha},\hat{c}_{\alpha},\hat{u}_{\alpha})=\frac{(v_{\alpha},c_{\alpha},u_{\alpha})}{\sqrt{RT_{0}}}$, $\hat{\rho}=\frac{\rho}{\rho_{0}}$, $\hat{T}=\frac{T}{T_{0}}$, $(\hat{\tau},\hat{t})=\frac{(\tau,t)}{L_{0}/\sqrt{RT_{0}}}$, and $\left(\hat{f},\hat{f}^{(0)},\hat{f}_{i},\hat{f}_{i}^{(0)}\right)=\frac{\left(f,f^{(0)},f_{i},f_{i}^{(0)}\right)}{\rho_{0}\left(RT_{0}\right)^{-3/2}}$. For simplicity, the symbol $\wedge$” associated with dimensionless variables will be omitted henceforth.
@@
In this section, the one-dimensional Riemann problem is employed to validate the model’s ability to describe large-scale structures. Specifically, Riemann problems, such as Sod and Lax shock tubes, involve mesoscale structures—shock waves, rarefaction waves, and contact discontinuities—where significant nonequilibrium effects arise. These problems also provide an opportunity to assess the model’s capability in describing small-scale structures and its ability to capture the multiscale nature of thermodynamic nonequilibrium effects.
@@ III.1.1 Einfeldt shock tube
$\left\{\begin{array}[]{l}\left(\rho,T,u_{x},u_{y},u_{z}\right)|_{L}=\left(1.0,1.0,0.75,0.0,0.0\right),\\ \left(\rho,T,u_{x},u_{y},u_{z}\right)|_{R}=\left(0.125,0.8,0.0,0.0,0.0\right).\end{array}\right.$ (44)
@@ III.1.2 Modified Sod shock tube
$\left\{\begin{array}[]{l}\left(\rho,T,u_{x},u_{y},u_{z}\right)|_{L}=\left(1.0,1.0,0.3,0.0,0.0\right),\\ \left(\rho,T,u_{x},u_{y},u_{z}\right)|_{R}=\left(0.125,0.25,-0.1,0.0,0.0\right).\end{array}\right.$ (49)
@@ III.1.3 Lax shock tube
$\left\{\begin{array}[]{l}\left(\rho,T,u_{x},u_{y},u_{z}\right)|_{L}=\left(0.445,7.928,0.698,0.0,0.0\right),\\ \left(\rho,T,u_{x},u_{y},u_{z}\right)|_{R}=\left(0.50,1.142,0.0,0.0,0.0\right).\end{array}\right.$ (50)
@@ III.1.4 Collision of two strong shocks
$\left\{\begin{array}[]{l}\left(\rho,T,u_{x},u_{y},u_{z}\right)|_{L}=\left(5.99924,76.8254,19.5975,0.0,0.0\right),\\ \left(\rho,T,u_{x},u_{y},u_{z}\right)|_{R}=\left(5.99242,7.69222,-6.19633,0.0,0.0\right).\end{array}\right.$ (51)
The numerical parameters for this problem are $\Delta x=\Delta y=\Delta z=4\times 10^{-3}$, $\Delta t=5\times 10^{-5}$, $\tau=5\times 10^{-5}$, $n=0$, $c=-13$, and $\eta_{0}=34$, while the remaining parameters remain unchanged. The numerical solutions for density, pressure, velocity, and temperature distributions closely align with the analytical solutions. The model handles strong shock waves with remarkable stability and accuracy, exhibiting minimal numerical dissipation [see Fig. 5]. It is worth noting that slight diffusion is observed in the steep shock wave regions. This diffusion is attributed to the DBM’s inherent dissipative mechanisms, such as viscous stress and heat conduction, enabling more effective representation of meso-scale structures like shock interfaces. This confirms the model’s robustness and stability in handling strong nonlinear compressible flows.
@@
This study analyzes six selected configurations, with their initial conditions provide in Table 1. The symbols $(P_{i},\rho_{i},u_{xi},u_{yi})$ represent the physical quantities in the $i$th quadrant, with $u_{z}=0$ across all quadrants. The simulation employs a grid of $N_{x}\times N_{y}\times N_{z}=200\times 200\times 4$, applying Dirichlet boundary conditions in the $x$ direction and periodic conditions in the $y$ and $z$ directions. The grid spacing and time step are set as $\Delta x=\Delta y=\Delta z=10^{-3}$, $\Delta t=10^{-4}$, and the relaxation time as $\tau=5\times 10^{-5}$.